# Spectral Dataset of Young Type Ib Supernovae and their Time-evolution

### v1

@@ Appendix A Observation Log
| SN (t days) | Date of Observation | Telescope/Instrument | Exposure (s) | Slit (arcsec) | Airmass |
| SN 2023ljf(-14.5) | 2023-06-23 09:29:05 | OGG 2m/FLOYDS | 3600 | 2″ | 1.42 |
| SN 2023ljf(-12.5) | 2023-06-25 08:16:46 | OGG 2m/FLOYDS | 3600 | 2″ | 1.42 |
| SN 2023ljf(-10.5) | 2023-06-27 08:21:33 | OGG 2m/FLOYDS | 2700 | 2″ | 1.19 |
| SN 2023ljf(-7.6) | 2023-06-30 06:25:18 | OGG 2m/FLOYDS | 2700 | 2″ | 1.04 |
| SN 2023ljf(-4.6) | 2023-07-03 07:23:36 | OGG 2m/FLOYDS | 2700 | 2″ | 1.12 |
| SN 2023ljf(19.4) | 2023-07-27 07:20:07 | OGG 2m/FLOYDS | 2700 | 2″ | 1.4 |
| SN 2023ljf(25.4) | 2023-08-02 05:57:19 | OGG 2m/FLOYDS | 2700 | 2″ | 1.12 |
| SN 2023ljf(28.4) | 2023-08-05 06:20:32 | OGG 2m/FLOYDS | 2700 | 2″ | 1.29 |
| SN 2023ljf(40.4) | 2023-08-17 05:59:48 | OGG 2m/FLOYDS | 2700 | 2″ | 1.41 |
| SN 2022nyo(-13.8)a | 2022-07-03 01:55:28 | SOAR 4.1m/GHTS BLUE | 1800 | 1″ | 1.43 |
| SN 2022nyo(-4.1)a | 2022-07-12 18:29:14 | SALT/RSS | 1533 | 1.5″ | 1.27 |
| SN 2022nyo(-1.5) | 2022-07-15 10:06:09 | COJ 2m/FLOYDS | 2700 | 2″ | 1.4 |
| SN 2022nyo(23.4) | 2022-08-09 08:30:45 | COJ 2m/FLOYDS | 2700 | 2″ | 1.42 |
| SN 2022nyo(33.4) | 2022-08-19 08:36:22 | COJ 2m/FLOYDS | 2700 | 2″ | 1.68 |
| SN 2021ukt(-12.2) | 2021-08-03 11:41:04 | OGG 2m/FLOYDS | 1200 | 2″ | 1.39 |
| SN 2021ukt(-9.1) | 2021-08-06 12:17:32 | OGG 2m/FLOYDS | 1200 | 2″ | 1.21 |
| SN 2021ukt(-6.2) | 2021-08-09 11:57:22 | OGG 2m/FLOYDS | 1200 | 2″ | 1.24 |
| SN 2021ukt(-0.1) | 2021-08-15 13:30:07 | OGG 2m/FLOYDS | 1200 | 2″ | 1.08 |
| SN 2021ukt(4.0) | 2021-08-19 14:38:17 | OGG 2m/FLOYDS | 1200 | 2″ | 1.16 |
| SN 2021ukt(12.0) | 2021-08-27 15:10:45 | COJ 2m/FLOYDS | 1200 | 2″ | 1.2 |
| SN 2021ukt(15.8) | 2021-08-31 10:25:41 | OGG 2m/FLOYDS | 1199 | 2″ | 1.25 |
| SN 2021ukt(19.8) | 2021-09-04 12:08:46 | OGG 2m/FLOYDS | 1200 | 2″ | 1.08 |
| SN 2021ukt(24.9) | 2021-09-09 12:43:45 | COJ 2m/FLOYDS | 2700 | 2″ | 1.47 |
| SN 2021ukt(33.8) | 2021-09-18 10:51:57 | OGG 2m/FLOYDS | 3600 | 2″ | 1.09 |
| SN 2021ukt(45.8) | 2021-09-30 09:59:09 | OGG 2m/FLOYDS | 3600 | 2″ | 1.09 |
| SN 2021ukt(54.8) | 2021-10-09 10:46:54 | COJ 2m/FLOYDS | 3600 | 2″ | 1.44 |
| SN 2021njk(-15.3) | 2021-05-26 10:39:13 | COJ 2m/FLOYDS | 2700 | 2″ | 1.20 |
| SN 2021njk(-11.3) | 2021-05-30 09:51:19 | COJ 2m/FLOYDS | 3600 | 2″ | 1.14 |
| SN 2021njk (-3.4) | 2021-06-07 08:07:57 | COJ 2m/FLOYDS | 1800 | 2″ | 1.02 |
| SN 2021njk(2.7) | 2021-06-13 10:05:47 | COJ 2m/FLOYDS | 1800 | 2″ | 1.3 |
| SN 2021njk(5.7) | 2021-06-16 10:20:45 | COJ 2m/FLOYDS | 1800 | 2″ | 1.42 |
| SN 2021njk(10.7) | 2021-06-21 09:43:39 | COJ 2m/FLOYDS | 1800 | 2″ | 1.34 |
| SN 2021njk(15.7) | 2021-06-26 09:47:11 | COJ 2m/FLOYDS | 2700 | 2″ | 1.51 |
| SN 2021njk(23.6) | 2021-07-04 09:31:19 | COJ 2m/FLOYDS | 2700 | 2″ | 1.61 |
| SN 2021hen(-8.2) | 2021-03-30 13:01:37 | OGG 2m/FLOYDS | 3600 | 2″ | 1.24 |
| SN 2021hen(-7.2) | 2021-03-31 11:53:45 | OGG 2m/FLOYDS | 3600 | 2″ | 1.08 |
| SN 2021hen(-1.2) | 2021-04-06 11:54:36 | OGG 2m/FLOYDS | 3600 | 2″ | 1.12 |
| SN 2021hen(1.9) | 2021-04-09 14:13:33 | COJ 2m/FLOYDS | 3600 | 2″ | 1.85 |
| SN 2021hen(5.6) | 2021-04-13 08:26:33 | OGG 2m/FLOYDS | 3600 | 2″ | 1.05 |
| SN 2021hen(11.8) | 2021-04-19 12:16:47 | OGG 2m/FLOYDS | 3600 | 2″ | 1.38 |
| SN 2020hvp(-13.2) | 2020-04-23 17:50:18 | COJ 2m/FLOYDS | 3600 | 2″ | 1.34 |
| SN 2020hvp(-11.2) | 2020-04-25 17:41:11 | COJ 2m/FLOYDS | 1800 | 2″ | 1.28 |
| SN 2020hvp(-9.6) | 2020-04-27 09:20:25 | OGG 2m/FLOYDS | 1800 | 2″ | 1.48 |
| SN 2020hvp(-8.3) | 2020-04-28 16:41:04 | COJ 2m/FLOYDS | 1200 | 2″ | 1.17 |
| SN 2020hvp(-7.4) | 2020-04-29 13:46:48 | OGG 2m/FLOYDS | 1200 | 2″ | 1.20 |
| SN 2020hvp(-5.4) | 2020-05-01 12:22:17 | OGG 2m/FLOYDS | 1200 | 2″ | 1.09 |
| SN 2020hvp(-4.4) | 2020-05-02 14:08:36 | OGG 2m/FLOYDS | 900 | 2″ | 1.30 |
| SN 2020hvp(-2.6) | 2020-05-04 08:53:23 | OGG 2m/FLOYDS | 900 | 2″ | 1.52 |
| SN 2020hvp(-1.4) | 2020-05-05 12:57:46 | OGG 2m/FLOYDS | 900 | 2″ | 1.15 |
| SN 2020hvp(-0.4) | 2020-05-06 14:08:49 | COJ 2m/FLOYDS | 900 | 2″ | 1.20 |
| SN 2020hvp(4.6) | 2020-05-11 13:07:53 | COJ 2m/FLOYDS | 900 | 2″ | 1.29 |
| SN 2020hvp(9.7) | 2020-05-16 16:25:52 | COJ 2m/FLOYDS | 900 | 2″ | 1.27 |
| SN 2020hvp(16.6) | 2020-05-23 12:49:18 | OGG 2m/FLOYDS | 1800 | 2″ | 1.35 |
| SN 2020hvp(27.6) | 2020-06-03 12:15:10 | OGG 2m/FLOYDS | 1800 | 2″ | 1.39 |
| SN 2020hvp(34.4) | 2020-06-10 07:46:35 | OGG 2m/FLOYDS | 2400 | 2″ | 1.17 |
| SN 2020hvp(51.6) | 2020-06-27 13:34:03 | COJ 2m/FLOYDS | 2700 | 2″ | 1.30 |
| SN 2019odp(-15.7) | 2019-08-24 12:44:36 | OGG 2m/FLOYDS | 3600 | 2″ | 1.14 |
| SN 2019odp(-11.6) | 2019-08-28 13:33:07 | OGG 2m/FLOYDS | 3599 | 2″ | 1.38 |
| SN 2019odp(-8.7) | 2019-08-31 13:35:03 | OGG 2m/FLOYDS | 3600 | 2″ | 1.46 |
| SN 2019odp(-5.6) | 2019-09-03 13:46:45 | COJ 2m/FLOYDS | 3600 | 2″ | 1.43 |
| SN 2019odp(-2.7) | 2019-09-06 11:08:12 | OGG 2m/FLOYDS | 2700 | 2″ | 1.04 |
| SN 2019odp(1.4) | 2019-09-10 13:47:02 | COJ 2m/FLOYDS | 2700 | 2″ | 1.43 |
| SN 2019odp(5.4) | 2019-09-14 14:48:18 | COJ 2m/FLOYDS | 2700 | 2″ | 1.58 |
| SN 2019odp(10.3) | 2019-09-19 12:10:09 | OGG 2m/FLOYDS | 2700 | 2″ | 1.35 |
| SN 2019odp(15.3) | 2019-09-24 11:23:11 | OGG 2m/FLOYDS | 2700 | 2″ | 1.23 |
| SN 2019odp(24.4) | 2019-10-03 13:29:53 | COJ 2m/FLOYDS | 2700 | 2″ | 1.57 |
| SN 2019odp(30.3) | 2019-10-09 12:35:08 | COJ 2m/FLOYDS | 2700 | 2″ | 1.48 |
| SN 2019odp(43.2) | 2019-10-22 09:12:29 | OGG 2m/FLOYDS | 3600 | 2″ | 1.2 |
| SN 2019odp(57.2) | 2019-11-05 10:23:11 | COJ 2m/FLOYDS | 3600 | 2″ | 1.46 |
| SN 2016bau(-15.9) | 2016-03-15 05:24:59 | OGG 2m/FLOYDS | 2700 | 2″ | 1.99 |
| SN 2016bau(-13.6) | 2016-03-17 13:01:07 | OGG 2m/FLOYDS | 2700 | 2″ | 1.56 |
| SN 2016bau(-10.6) | 2016-03-20 12:47:21 | OGG 2m/FLOYDS | 2700 | 2″ | 1.55 |
| SN 2016bau(-6.7) | 2016-03-24 12:04:09 | OGG 2m/FLOYDS | 2700 | 2″ | 1.44 |
| SN 2016bau(-1.7) | 2016-03-29 10:25:44 | OGG 2m/FLOYDS | 2700 | 2″ | 1.25 |
| SN 2016bau(5.1) | 2016-04-05 06:56:20 | OGG 2m/FLOYDS | 2700 | 2″ | 1.25 |
| SN 2016bau(9.2) | 2016-04-09 08:26:46 | OGG 2m/FLOYDS | 2700 | 2″ | 1.19 |
| SN 2016bau(10.3) | 2016-04-10 11:53:54 | OGG 2m/FLOYDS | 2700 | 2″ | 1.7 |
| SN 2016bau(15.2) | 2016-04-15 08:53:05 | OGG 2m/FLOYDS | 2700 | 2″ | 1.22 |
| SN 2016bau(24.2) | 2016-04-24 08:37:56 | OGG 2m/FLOYDS | 2700 | 2″ | 1.24 |
| SN 2016bau(32.2) | 2016-05-02 09:18:02 | OGG 2m/FLOYDS | 2700 | 2″ | 1.4 |
| SN 2016bau(42.2) | 2016-05-12 09:28:37 | OGG 2m/FLOYDS | 2700 | 2″ | 1.59 |
| SN 2016bau(50.2) | 2016-05-20 08:55:44 | OGG 2m/FLOYDS | 2700 | 2″ | 1.59 |
Note: Host galaxy HII region emission lines are removed from these spectra. Figure 1 shows the spectra without the galaxy emission lines.

### v2

@@ Appendix B Observation log
| SN (t days) | type | Date of Observation | Telescope/Instrument | Exposure (s) | Slit () | Airmass |
| SN 2023ljf(14.5) | Ib | 2023-06-23 09:29:05 | OGG 2m/FLOYDS | 3600 | 2″ | 1.42 |
| SN 2023ljf(12.5) | Ib | 2023-06-25 08:16:46 | OGG 2m/FLOYDS | 3600 | 2″ | 1.42 |
| SN 2023ljf(10.5) | Ib | 2023-06-27 08:21:33 | OGG 2m/FLOYDS | 2700 | 2″ | 1.19 |
| SN 2023ljf(7.6) | Ib | 2023-06-30 06:25:18 | OGG 2m/FLOYDS | 2700 | 2″ | 1.04 |
| SN 2023ljf(4.6) | Ib | 2023-07-03 07:23:36 | OGG 2m/FLOYDS | 2700 | 2″ | 1.12 |
| SN 2023ljf(19.4) | Ib | 2023-07-27 07:20:07 | OGG 2m/FLOYDS | 2700 | 2″ | 1.4 |
| SN 2023ljf(25.4) | Ib | 2023-08-02 05:57:19 | OGG 2m/FLOYDS | 2700 | 2″ | 1.12 |
| SN 2023ljf(28.4) | Ib | 2023-08-05 06:20:32 | OGG 2m/FLOYDS | 2700 | 2″ | 1.29 |
| SN 2023ljf(40.4) | Ib | 2023-08-17 05:59:48 | OGG 2m/FLOYDS | 2700 | 2″ | 1.41 |
| SN 2022nyo(13.8)a | IIb | 2022-07-03 01:55:28 | SOAR 4.1m/GHTS BLUE | 1800 | 1″ | 1.43 |
| SN 2022nyo(4.1)a | Ib | 2022-07-12 18:29:14 | SALT/RSS | 1533 | 1.5″ | 1.27 |
| SN 2022nyo(1.5) | Ib | 2022-07-15 10:06:09 | COJ 2m/FLOYDS | 2700 | 2″ | 1.4 |
| SN 2022nyo(23.4) | Ib | 2022-08-09 08:30:45 | COJ 2m/FLOYDS | 2700 | 2″ | 1.42 |
| SN 2022nyo(33.4) | Ib | 2022-08-19 08:36:22 | COJ 2m/FLOYDS | 2700 | 2″ | 1.68 |
| SN 2021ukt(12.2) | IIn | 2021-08-03 11:41:04 | OGG 2m/FLOYDS | 1200 | 2″ | 1.39 |
| SN 2021ukt(9.1) | Ib | 2021-08-06 12:17:32 | OGG 2m/FLOYDS | 1200 | 2″ | 1.21 |
| SN 2021ukt(6.2) | Ib | 2021-08-09 11:57:22 | OGG 2m/FLOYDS | 1200 | 2″ | 1.24 |
| SN 2021ukt(0.1) | Ib | 2021-08-15 13:30:07 | OGG 2m/FLOYDS | 1200 | 2″ | 1.08 |
| SN 2021ukt(4.0) | Ib | 2021-08-19 14:38:17 | OGG 2m/FLOYDS | 1200 | 2″ | 1.16 |
| SN 2021ukt(12.0) | Ib | 2021-08-27 15:10:45 | COJ 2m/FLOYDS | 1200 | 2″ | 1.2 |
| SN 2021ukt(15.8) | Ib | 2021-08-31 10:25:41 | OGG 2m/FLOYDS | 1199 | 2″ | 1.25 |
| SN 2021ukt(19.8) | Ib | 2021-09-04 12:08:46 | OGG 2m/FLOYDS | 1200 | 2″ | 1.08 |
| SN 2021ukt(24.9) | Ib | 2021-09-09 12:43:45 | COJ 2m/FLOYDS | 2700 | 2″ | 1.47 |
| SN 2021ukt(33.8) | Ib | 2021-09-18 10:51:57 | OGG 2m/FLOYDS | 3600 | 2″ | 1.09 |
| SN 2021ukt(45.8) | Ib | 2021-09-30 09:59:09 | OGG 2m/FLOYDS | 3600 | 2″ | 1.09 |
| SN 2021ukt(54.8) | Ib | 2021-10-09 10:46:54 | COJ 2m/FLOYDS | 3600 | 2″ | 1.44 |
| SN 2021njk(15.3) | Ib | 2021-05-26 10:39:13 | COJ 2m/FLOYDS | 2700 | 2″ | 1.20 |
| SN 2021njk(11.3) | Ib | 2021-05-30 09:51:19 | COJ 2m/FLOYDS | 3600 | 2″ | 1.14 |
| SN 2021njk (3.4) | Ib | 2021-06-07 08:07:57 | COJ 2m/FLOYDS | 1800 | 2″ | 1.02 |
| SN 2021njk(2.7) | Ib | 2021-06-13 10:05:47 | COJ 2m/FLOYDS | 1800 | 2″ | 1.3 |
| SN 2021njk(5.7) | Ib | 2021-06-16 10:20:45 | COJ 2m/FLOYDS | 1800 | 2″ | 1.42 |
| SN 2021njk(10.7) | Ib | 2021-06-21 09:43:39 | COJ 2m/FLOYDS | 1800 | 2″ | 1.34 |
| SN 2021njk(15.7) | Ib | 2021-06-26 09:47:11 | COJ 2m/FLOYDS | 2700 | 2″ | 1.51 |
| SN 2021njk(23.6) | Ib | 2021-07-04 09:31:19 | COJ 2m/FLOYDS | 2700 | 2″ | 1.61 |
| SN 2021hen(8.2) | Ib | 2021-03-30 13:01:37 | OGG 2m/FLOYDS | 3600 | 2″ | 1.24 |
| SN 2021hen(7.2) | Ib | 2021-03-31 11:53:45 | OGG 2m/FLOYDS | 3600 | 2″ | 1.08 |
| SN 2021hen(1.2) | Ib | 2021-04-06 11:54:36 | OGG 2m/FLOYDS | 3600 | 2″ | 1.12 |
| SN 2021hen(1.9) | Ib | 2021-04-09 14:13:33 | COJ 2m/FLOYDS | 3600 | 2″ | 1.85 |
| SN 2021hen(5.6) | Ib | 2021-04-13 08:26:33 | OGG 2m/FLOYDS | 3600 | 2″ | 1.05 |
| SN 2021hen(11.8) | Ib | 2021-04-19 12:16:47 | OGG 2m/FLOYDS | 3600 | 2″ | 1.38 |
| SN 2020hvp(13.2) | Ib | 2020-04-23 17:50:18 | COJ 2m/FLOYDS | 3600 | 2″ | 1.34 |
| SN 2020hvp(11.2) | Ib | 2020-04-25 17:41:11 | COJ 2m/FLOYDS | 1800 | 2″ | 1.28 |
| SN 2020hvp(9.6) | Ib | 2020-04-27 09:20:25 | OGG 2m/FLOYDS | 1800 | 2″ | 1.48 |
| SN 2020hvp(8.3) | Ib | 2020-04-28 16:41:04 | COJ 2m/FLOYDS | 1200 | 2″ | 1.17 |
| SN 2020hvp(7.4) | Ib | 2020-04-29 13:46:48 | OGG 2m/FLOYDS | 1200 | 2″ | 1.20 |
| SN 2020hvp(5.4) | Ib | 2020-05-01 12:22:17 | OGG 2m/FLOYDS | 1200 | 2″ | 1.09 |
| SN 2020hvp(4.4) | Ib | 2020-05-02 14:08:36 | OGG 2m/FLOYDS | 900 | 2″ | 1.30 |
| SN 2020hvp(2.6) | Ib | 2020-05-04 08:53:23 | OGG 2m/FLOYDS | 900 | 2″ | 1.52 |
| SN 2020hvp(1.4) | Ib | 2020-05-05 12:57:46 | OGG 2m/FLOYDS | 900 | 2″ | 1.15 |
| SN 2020hvp(0.4) | Ib | 2020-05-06 14:08:49 | COJ 2m/FLOYDS | 900 | 2″ | 1.20 |
| SN 2020hvp(4.6) | Ib | 2020-05-11 13:07:53 | COJ 2m/FLOYDS | 900 | 2″ | 1.29 |
| SN 2020hvp(9.7) | Ib | 2020-05-16 16:25:52 | COJ 2m/FLOYDS | 900 | 2″ | 1.27 |
| SN 2020hvp(16.6) | Ib | 2020-05-23 12:49:18 | OGG 2m/FLOYDS | 1800 | 2″ | 1.35 |
| SN 2020hvp(27.6) | Ib | 2020-06-03 12:15:10 | OGG 2m/FLOYDS | 1800 | 2″ | 1.39 |
| SN 2020hvp(34.4) | Ib | 2020-06-10 07:46:35 | OGG 2m/FLOYDS | 2400 | 2″ | 1.17 |
| SN 2020hvp(51.6) | Ib | 2020-06-27 13:34:03 | COJ 2m/FLOYDS | 2700 | 2″ | 1.30 |
| SN 2019odp(15.7) | Ic-bl | 2019-08-24 12:44:36 | OGG 2m/FLOYDS | 3600 | 2″ | 1.14 |
| SN 2019odp(11.6) | Ic-bl | 2019-08-28 13:33:07 | OGG 2m/FLOYDS | 3599 | 2″ | 1.38 |
| SN 2019odp(8.7) | Ic-bl | 2019-08-31 13:35:03 | OGG 2m/FLOYDS | 3600 | 2″ | 1.46 |
| SN 2019odp(5.6) | Ic-bl | 2019-09-03 13:46:45 | COJ 2m/FLOYDS | 3600 | 2″ | 1.43 |
| SN 2019odp(2.7) | Ic-bl | 2019-09-06 11:08:12 | OGG 2m/FLOYDS | 2700 | 2″ | 1.04 |
| SN 2019odp(1.4) | Ib | 2019-09-10 13:47:02 | COJ 2m/FLOYDS | 2700 | 2″ | 1.43 |
| SN 2019odp(5.4) | Ib | 2019-09-14 14:48:18 | COJ 2m/FLOYDS | 2700 | 2″ | 1.58 |
| SN 2019odp(10.3) | Ib | 2019-09-19 12:10:09 | OGG 2m/FLOYDS | 2700 | 2″ | 1.35 |
| SN 2019odp(15.3) | Ib | 2019-09-24 11:23:11 | OGG 2m/FLOYDS | 2700 | 2″ | 1.23 |
| SN 2019odp(24.4) | Ib | 2019-10-03 13:29:53 | COJ 2m/FLOYDS | 2700 | 2″ | 1.57 |
| SN 2019odp(30.3) | Ib | 2019-10-09 12:35:08 | COJ 2m/FLOYDS | 2700 | 2″ | 1.48 |
| SN 2019odp(43.2) | Ib | 2019-10-22 09:12:29 | OGG 2m/FLOYDS | 3600 | 2″ | 1.2 |
| SN 2019odp(57.2) | Ib | 2019-11-05 10:23:11 | COJ 2m/FLOYDS | 3600 | 2″ | 1.46 |
| SN 2016bau(15.9) | Ib | 2016-03-15 05:24:59 | OGG 2m/FLOYDS | 2700 | 2″ | 1.99 |
| SN 2016bau(13.6) | Ib | 2016-03-17 13:01:07 | OGG 2m/FLOYDS | 2700 | 2″ | 1.56 |
| SN 2016bau(10.6) | Ib | 2016-03-20 12:47:21 | OGG 2m/FLOYDS | 2700 | 2″ | 1.55 |
| SN 2016bau(6.7) | Ib | 2016-03-24 12:04:09 | OGG 2m/FLOYDS | 2700 | 2″ | 1.44 |
| SN 2016bau(1.7) | Ib | 2016-03-29 10:25:44 | OGG 2m/FLOYDS | 2700 | 2″ | 1.25 |
| SN 2016bau(5.1) | Ib | 2016-04-05 06:56:20 | OGG 2m/FLOYDS | 2700 | 2″ | 1.25 |
| SN 2016bau(9.2) | Ib | 2016-04-09 08:26:46 | OGG 2m/FLOYDS | 2700 | 2″ | 1.19 |
| SN 2016bau(10.3) | Ib | 2016-04-10 11:53:54 | OGG 2m/FLOYDS | 2700 | 2″ | 1.7 |
| SN 2016bau(15.2) | Ib | 2016-04-15 08:53:05 | OGG 2m/FLOYDS | 2700 | 2″ | 1.22 |
| SN 2016bau(24.2) | Ib | 2016-04-24 08:37:56 | OGG 2m/FLOYDS | 2700 | 2″ | 1.24 |
| SN 2016bau(32.2) | Ib | 2016-05-02 09:18:02 | OGG 2m/FLOYDS | 2700 | 2″ | 1.4 |
| SN 2016bau(42.2) | Ib | 2016-05-12 09:28:37 | OGG 2m/FLOYDS | 2700 | 2″ | 1.59 |
| SN 2016bau(50.2) | Ib | 2016-05-20 08:55:44 | OGG 2m/FLOYDS | 2700 | 2″ | 1.59 |